I have been editing and reloading the c1ioo model last two days. I have restarted the frame builder several times. After one of the restarts on Sunday evening the fb started having problems which initially showed up as dtt reporting synchronization error. This morning Kiwamu and I tried to restart the fb again and it stopped working all together. We called Joe and he fixed the fb problem by fixing the time stamps (Joe will add details to describe the fix when he sees this elog).
The following changes were made to c1ioo model:
- The angular dither lockins were added for each optics to do the beam spot centering on MC mirrors. The MCL signal is demodulated digitally at 3 pitch and 3 yaw frequencies. (The MCL signal was reconnected to the first input of the ADC interface board).
- The outputs of the lockins go through the sensing matrix, DOF filters, and control matrix to the MC1,2,3 SUS-MC1(2,3)_ASCPIT(YAW) filter inputs where they sum with dither signals (CLOCK output of the oscillators).
- The MCL_TEST_FILT was removed
The arm cavity dither alignment (c1ass) status:
- The demodulated signals were minimized by moving the ETMX/ITMX optic biases and simultaneously keeping the arm buildup (TRX) high by using the BS and PZT2. The minimization of the TRX demodulated signals has not been successful for some reason.
- The next step is to close the servo loops REFL11I demodulated signals -> TMs and TRX demodulated signals -> combination of BS and PZTs.
The MC dither alignment (c1ioo) status:
- The demodulated signals were obtained and sensing matrix (MCs -> lockin outputs) was measured for pitch dof.
- The inversion of the matrix is in progress.
- The additional c1ass and c1ioo medm screens and up and down scripts are being made.